# Improving Graph Neural Networks by Learning Continuous Edge Directions

### v1

@@ A.2.1 Data generation and training setup for the directed flow triangular lattice graph
Model # layers # hidden lr self-loop sc # attn. heads attn. type enc. type K aggr. # enc. layers # dec. layer self-feature lr $\Theta$ lw $\Theta$
GCN 2 16 1e-3 - - - - - - - - - - - -
GAT 4 32 1e-3 - True 4 - - - - - - - - -
GraphGPS 4 64 1e-3 0 - 4 multihead eigenvector (dim=5) - - - - - - -
@@
DRew 3 64 5e-3 - - - - - - weight - - - - -
FLODE 4 64 1e-2 0 - - - - - - 1 3 - - -
CoED 4 64 1e-3 - - - - - - - - - True 1e-3 False
@@ A.2.2 Data generation and Training setup for the GRN dynamics experiment
Model # layers # hidden lr self-loop sc # attn. heads attn. type enc. type K aggr. # enc. layers # dec. layer self-feature lr $\Theta$ lw $\Theta$
GCN 3 32 5e-4 - - - - - - - - - - - -
GAT 5 32 2e-3 - True 8 - - - - - - - - -
GraphGPS 5 32 5e-3 0 - 4 multihead eigenvector (dim=10) - - - - - - -
MagNet 5 32 5e-3 1 - - - - 2 - - - - - -
DRew 3 32 5e-4 - - - - - - weight - - - - -
FLODE 5 32 5e-3 1 - - - - - - 1 1 - - -
CoED 5 32 1e-3 - - - - - - - - - True 1e-2 True
@@ A.3.3 Training setup for power grid experiments
    For each bus node, integrate the features of its subnodes into its own features via GraphConv (i.e., $\mathbf{W}_{1}\mathbf{x}_{i}+\mathbf{W}_{2}\sum_{j\in\mathcal{N}(i)}e_{j,i}\cdot\mathbf{x}_{j}$ ).
@@ D.2 Fuzzy Laplacian Matrix
$(\mathbf{L}_{F})_{ij}=\begin{cases}0&\text{if }A_{ij}=A_{ji}=0\\ e^{i\theta_{ij}}&\text{otherwise}\end{cases}$ (6)
@@
In other words, the real part of $e^{i\theta_{ij}}$ is equal to the corresponding adjacency matrix entry $A_{ij}$. We require that $20\leq\theta_{ij}\leq\pi/$ 2. It follows that $\theta_{ji}=\pi/2-\theta_{ij}$.
@@
$\mathbf{L}_{F}=i\mathbf{L}_{F}^{*}$
To confirm this, note that $e^{-i\theta_{ij}}=\cos(-\theta_{ij})+i\sin(-\theta_{ij})=\cos(\theta_{ij})-i\sin(\theta_{ij})$. Therefore, $(\mathbf{L}_{F})_{ji}=\sin(\theta_{ij})+i\cos(\theta_{ij})=ie^{-i\theta_{ij}}$, and thus $\mathbf{L}_{F}=i\mathbf{L}_{F}^{*}$.
@@
$\mathbf{L}_{F}=\begin{pmatrix}0&\cdots&\cdots\\ \vdots&\ddots&e^{i\theta_{ij}}\\ \vdots&ie^{-i\theta_{ij}}&\ddots\\ \end{pmatrix}$ (7)
Here, $e^{i\theta_{ij}}$ and $ie^{-i\theta_{ij}}$ are sample off-diagonal elements corresponding to the edge $(v_{i},v_{j})$ in the graph.
@@ D.3 Properties of Fuzzy Laplacian Matrix $\mathbf{L}_{F}$
In this section, we will show that the Fuzzy Laplacian matrix $\mathbf{L}_{F}$ has eigenvalues of the form $a+ia$, where $a\in\mathbb{R}$, and orthogonal eigenvectors.
D.3.1 Eigenvalues of the Form $a+ia$
$\mathbf{L}_{F}$ has eigenvalues of the form $a+ia$ with $a\in\mathbb{R}$.
@@ D.3.1 Eigenvalues of the Form $a+ia$
$\Rightarrow\ \ -i\mathbf{w}^{*}\mathbf{L}_{F}\mathbf{w}=\lambda^{*}\mathbf{w}^{*}\mathbf{w}\ \ \Rightarrow\ \ -i\lambda\mathbf{w}^{*}\mathbf{w}=\lambda^{*}\mathbf{w}^{*}\mathbf{w}\ \ \Rightarrow\ \ -i\lambda=\lambda^{*}$
where we used $\mathbf{L}_{F}^{*}=-i\mathbf{L}_{F}$ to go to the second line. The last identity holds only when $\lambda=a+ia$ where $a$ is a real number.
@@ D.3.2 Orthogonal Eigenvectors
$\mathbf{v}^{*}\mathbf{L}_{F}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ i\mathbf{v}^{*}\mathbf{L}_{F}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ i\lambda_{2}^{*}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ \lambda_{2}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}$
where we used $\lambda=i\lambda^{*}$ derived above for the last step. Therefore, $\lambda_{2}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}$. Since $\lambda_{1}\neq\lambda_{2}$, it must be that $\mathbf{v}^{*}\mathbf{w}=0$, i.e., $\mathbf{w}$ and $\mathbf{v}$ are orthogonal.
@@ E.2 Weisfeiler-Leman test for isomorphism of directed graphs with fuzzy edges
WL test is a vertex refinement algorithm that assigns starting features to each node of the graph. The algorithm then aggregates all the features of each node’s neighbors and hashes the aggregated labels alongside the node’s own label into a unique new label. At each iteration of the algorithm, the list of labels are compared across the two graphs. If they are different, the two graphs are not isomorphic. If the labels are no longer updated at each iteration, the two graphs can potentially be isomorphic.
We extend the WL test to directed graphs with fuzzy edges in two ways. Importantly, for all the proofs that follow, we will assume that the graph has a finite number of nodes.
@@
The proof of above statement is a trivial extension of the proof of Theorem 3 in (Xu et al., 2018). Briefly, the multiset of the features of the neighbors of node $i$ can be converted from a multiset of tuples of the form $\{\!\!\{(\mu_{ij},h^{(k-1)}_{j}):j\in N(i)\}\!\!\}$ to a multiset of augmented features of the neighbors $\{\!\!\{\tilde{h}^{(k-1)}_{j}:j\in N(i)\}\!\!\}$ where $\tilde{h}^{(k-1)}_{j}=(\mu_{ij},h^{(k-1)}_{j})$. Because $\mu_{ij}$ form a set of finite cardinality, the augmented feature space $\tilde{h}^{(k-1)}_{j}$ is still a countable set and the same proof as in (Xu et al., 2018) can be applied.
@@
Note that $F(i)=\sum_{j\in N(i)}\mu_{ij}f(h^{(k-1)}_{j})$ and $G(i)=\sum_{j\in N(i)}\mu_{ji}f(h^{(k-1)}_{j})$ are only injective up to the sum of the incoming and outgoing weights for a given type of neighbor (see Figure E.2 for an example). This is because the weak form of the WL test only accounts for the sum of the weights of outgoing and incoming edges to neighboring nodes of a given color. This is different from previous work that dealt with undirected graphs (Xu et al., 2018) or graphs with directed but unweighted edges (Rossi et al., 2024).
@@
Because our graphs have a finite number of nodes, $X$ is bounded. For graphs without directed edges or weighted edges, it can be shown (see Lemma 5 of (Xu et al., 2018)) that there always exists an injective function $f$ such that $F(x)=\sum_{x\in X}f(x)$ is injective for all multisets $X$, i.e. if $F(X)=F(Y)$ for two multisets $X$ and $Y$ then $X=Y$. This result can be easily extended to regular directed graphs without weighted edges by considering the neighbors with incoming edges separately from the neighbors with outgoing edges (Rossi et al., 2024). For such graphs, a function $f$ exists such that the tuple $(\sum_{x\in X_{\rightarrow}}f(x),\sum_{x\in X_{\leftarrow}}f(x))$ is only the same for two nodes if they have identical neighborhoods. $X_{\rightarrow}$ and $X_{\leftarrow}$ denotes the multiset of features of the neighboring nodes that are connected to $i$ using an outgoing edge and incoming edge respectively.
@@ F.1 Magnetic Laplacian
$\mathbf{H}^{(q)}=S\odot\Theta({q}).$
$\mathbf{H}^{(q)}$ has some useful properties in capturing the directionality of the edges of the graph. For example, for $q=1/4$, if there is an edge connecting $j$ to $k$ but no edge connecting $k$ to $j$ then $\mathbf{H}^{(q)}_{jk}=i/2$ and $\mathbf{H}^{(q)}_{kj}=-i/2$. Although this encoding of edge direction is useful, the fact that both incoming and outgoing edges are purely imaginary (only different up to a sign) means that in general it is impossible to distinguish features aggregated from neighbors that are connected to a node through outgoing edges from those connected through incoming edges. The fuzzy Laplacian $\textbf{L}_{F}$, however, trivially distinguishes features from outgoing neighbors from those of incoming neighbors by keeping one set real and the other imaginary. We will expand on the implication of this for the expressivity of graph neural networks constructed using these two approaches below.
@@ F.2 Limitations of the magnetic Laplacian
Proof.. To prove this statement, we will show that the MLP can learn a linear combination of its input that would it make it equivalent to the graph neural network defined in Eq. 11. This is because when $q=1/8$, the real and imaginary parts of the aggregated features of the neighboring nodes are linear combinations of the features of the outgoing and incoming neighbors.
@@
To extend the magnetic Laplacian to directed graphs with fuzzy edges, we can assign a different value for $q$ (Eq. 15) to each edge. To keep the notation comparable to that of the fuzzy Laplacian, we instead assign a separate angle $\theta_{ij}$ to each edge and define the magnetic Laplacian as, $(\mathbf{L}_{M})_{ij}=e^{i\theta_{ij}}$ for $j>i$ if there is an edge between the nodes $i$ and $j$, $S_{ij}\neq 0$. To ensure that magnetic Laplacian remains Hermitian, $\mathbf{L}_{M}^{*}=\mathbf{L}_{M}$, we require that $(\mathbf{L}_{M})_{ji}=e^{-i\theta_{ij}}$, which defines the $\theta_{ij}$ values for $j<i$. It follows that $\theta_{ji}=-\theta_{ij}$.
$\mathbf{L}_{M}=\begin{pmatrix}0&\cdots&\cdots\\ \vdots&\ddots&e^{i\theta_{ij}}\\ \vdots&e^{-i\theta_{ij}}&\ddots\\ \end{pmatrix}$
@@
Lemma. For directed graphs with fuzzy edges, the graph neural network constructed using the magnetic Laplacian $\mathbf{L}_{M}$, Eq. 16, is not as expressive as the weak form of the WL graph isomorphism test, and therefore not as expressive as the graph neural network defined using the fuzzy Laplacian $\mathbf{L}_{F}$, Eq. 11.
Proof. The Fuzzy Laplacian matrix conveniently captures the outgoing and incoming weights of the edges from one node to another, $(\mathbf{L}_{F})_{ij}=\cos(\theta_{ij})+i\sin(\theta_{ij})$. Namely, $\Re(\mathbf{L}_{F})_{ij}=\cos(\theta_{ij})$ is the outgoing weight $\mu_{ij}$ of node $i$ to node $j$ and $\Im(\mathbf{L}_{F})_{ij}=\sin(\theta_{ij})$ is in the incoming weight from node $j$ to node $i$, $\mu_{ji}$. Importantly, the outgoing weight from node $i$ to node $j$, $\mu_{ij}$, is the same as the incoming weight from node $i$ to node $j$. Similarly, the incoming weight from node $j$ to node $i$, $\mu_{ji}$ is the same as the outgoing weight from node $j$ to node $i$. These relationships are directly captured in the fuzzy Laplacian because $(\mathbf{L}_{F})_{ji}=\sin(\theta_{ij})+i\cos(\theta_{ij})$.
@@
From our definition above, $(\mathbf{L}_{M})_{ij}=\cos(\theta_{ij})+i\sin(\theta_{ij})$ and $(\mathbf{L}_{M})_{ji}=\cos(\theta_{ij})-i\sin(\theta_{ij})$. To relate $\theta_{ij}$ to $\mu_{ij}$ we note that the ratio of outgoing and incoming weights for node $i$ is $\mu_{ij}/\mu_{ji}$ and for node $j$ is $\mu_{ji}/\mu_{ij}$. Because these two quantities are reciprocals of each other, we can define

### v2

@@ A.2.1 Data generation and training setup for the directed flow triangular lattice graph
Model # layers # hidden lr self-loop sc # attn. heads attn. type enc. type K $\alpha$ aggr. # enc. layers # dec. layer self-feature lr $\Theta$ lw $\Theta$
GCN 2 16 1e-3 - - - - - - - - - - - - -
GAT 4 32 1e-3 - True 4 - - - - - - - - - -
GraphGPS 4 64 1e-3 0 - 4 multihead eigenvector (dim=5) - - - - - - - -
@@
Chung 4 64 1e-3 0 - - - - - 0.5 - - - - - -
DRew 3 64 5e-3 - - - - - - - weight - - - - -
FLODE 4 64 1e-2 0 - - - - - - - 1 3 - - -
CoED 4 64 1e-3 - - - - - - - - - - True 1e-3 False
@@ A.2.2 Data generation and Training setup for the GRN dynamics experiment
Model # layers # hidden lr self-loop sc # attn. heads attn. type enc. type K $\alpha$ aggr. # enc. layers # dec. layer self-feature lr $\Theta$ lw $\Theta$
GCN 3 32 5e-4 - - - - - - - - - - - - -
GAT 5 32 2e-3 - True 8 - - - - - - - - - -
GraphGPS 5 32 5e-3 0 - 4 multihead eigenvector (dim=10) - - - - - - - -
MagNet 5 32 5e-3 1 - - - - 2 - - - - - - -
Chung 5 32 1e-3 - - - - - - 0.5 - - - True
DRew 3 32 5e-4 - - - - - - - weight - - - - -
FLODE 5 32 5e-3 1 - - - - - - - 1 1 - - -
CoED 5 32 1e-3 - - - - - - - - - - True 1e-2 True
@@ A.3.3 Training setup for power grid experiments
    For each bus node, integrate the features of its subnodes into its own features via GraphConv (i.e., $\mathbf{W}_{1}\mathbf{x}_{i}+\mathbf{W}_{2}\sum_{j\in\mathcal{N}(i)}e_{ji}\cdot\mathbf{x}_{j}$ ).
@@ D.2 Fuzzy Laplacian Matrix
$(\mathbf{L}_{F})_{ij}=\begin{cases}0&\text{if }A_{ij}=A_{ji}=0\\ e^{\mathrm{i}\theta_{ij}}&\text{otherwise}\end{cases}$ (6)
@@
In other words, the real part of $e^{\mathrm{i}\theta_{ij}}$ is equal to the corresponding adjacency matrix entry $A_{ij}$. We require that $20\leq\theta_{ij}\leq\pi/$ 2. It follows that $\theta_{ji}=\pi/2-\theta_{ij}$.
@@
$\mathbf{L}_{F}=\mathrm{i}\mathbf{L}_{F}^{*}$
To confirm this, note that $e^{-\mathrm{i}\theta_{ij}}=\cos(-\theta_{ij})+\mathrm{i}\sin(-\theta_{ij})=\cos(\theta_{ij})-\mathrm{i}\sin(\theta_{ij})$. Therefore, $(\mathbf{L}_{F})_{ji}=\sin(\theta_{ij})+\mathrm{i}\cos(\theta_{ij})=\mathrm{i}e^{-\mathrm{i}\theta_{ij}}$, and thus $\mathbf{L}_{F}=\mathrm{i}\mathbf{L}_{F}^{*}$.
@@
$\mathbf{L}_{F}=\begin{pmatrix}0&\cdots&\cdots\\ \vdots&\ddots&e^{\mathrm{i}\theta_{ij}}\\ \vdots&\mathrm{i}e^{-\mathrm{i}\theta_{ij}}&\ddots\\ \end{pmatrix}$ (7)
Here, $e^{\mathrm{i}\theta_{ij}}$ and $\mathrm{i}e^{-\mathrm{i}\theta_{ij}}$ are sample off-diagonal elements corresponding to the edge $(v_{i},v_{j})$ in the graph.
@@ D.3 Properties of Fuzzy Laplacian Matrix $\mathbf{L}_{F}$
In this section, we will show that the Fuzzy Laplacian matrix $\mathbf{L}_{F}$ has eigenvalues of the form $a+\mathrm{i}a$, where $a\in\mathbb{R}$, and orthogonal eigenvectors.
D.3.1 Eigenvalues of the Form $a+\mathrm{i}a$
$\mathbf{L}_{F}$ has eigenvalues of the form $a+\mathrm{i}a$ with $a\in\mathbb{R}$.
@@ D.3.1 Eigenvalues of the Form $a+\mathrm{i}a$
$\Rightarrow\ \ -\mathrm{i}\mathbf{w}^{*}\mathbf{L}_{F}\mathbf{w}=\lambda^{*}\mathbf{w}^{*}\mathbf{w}\ \ \Rightarrow\ \ -\mathrm{i}\lambda\mathbf{w}^{*}\mathbf{w}=\lambda^{*}\mathbf{w}^{*}\mathbf{w}\ \ \Rightarrow\ \ -\mathrm{i}\lambda=\lambda^{*}$
where we used $\mathbf{L}_{F}^{*}=-\mathrm{i}\mathbf{L}_{F}$ to go to the second line. The last identity holds only when $\lambda=a+\mathrm{i}a$ where $a$ is a real number.
@@ D.3.2 Orthogonal Eigenvectors
$\mathbf{v}^{*}\mathbf{L}_{F}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ \mathrm{i}\mathbf{v}^{*}\mathbf{L}_{F}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ \mathrm{i}\lambda_{2}^{*}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}\ \ \Rightarrow\ \ \lambda_{2}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}$
where we used $\lambda=\mathrm{i}\lambda^{*}$ derived above for the last step. Therefore, $\lambda_{2}\mathbf{v}^{*}\mathbf{w}=\lambda_{1}\mathbf{v}^{*}\mathbf{w}$. Since $\lambda_{1}\neq\lambda_{2}$, it must be that $\mathbf{v}^{*}\mathbf{w}=0$, i.e., $\mathbf{w}$ and $\mathbf{v}$ are orthogonal.
@@ E.2 Weisfeiler-Leman test for isomorphism of directed graphs with fuzzy edges
WL test is a vertex refinement algorithm that assigns starting features to each node of the graph. The algorithm then aggregates all the features of each node’s neighbors and hashes the aggregated labels alongside the node’s own label into a unique new label. At each iteration of the algorithm, the list of labels is compared across the two graphs. If the labels are different, the two graphs are not isomorphic. If the labels are no longer updated at each iteration, the two graphs can potentially be isomorphic.
The Weisfeiler-Lehman (WL) graph isomorphism test treats all edges emanating from a node equivalently, as long as they connect to neighbors with the same label. Neighbors are distinguished solely by their labels, not by the attributes of the edges that connect them. When extending the WL test to directed graphs with fuzzy edges, we must account for the fact that fuzzy edges disrupt this uniform treatment of edges. We introduce the following two extensions of the WL test to handle directed graphs with fuzzy edges.
  • Strong form: A node can distinguish its neighbors not only by their labels but also by the weights of the edges connecting them. This allows the node to differentiate between neighbors of the same label based on edge attributes.
  • Weak form: A node cannot differentiate between neighbors of the same label based on edge weights. Instead, the only weight-related information a node can use is the sum of the edge weights connecting it to all neighbors with a given label.
These extensions adapt the WL test to account for the additional complexity introduced by fuzzy edges in directed graphs. Importantly, for all the proofs that follow, we will assume that the graph has a finite number of nodes.
@@
The proof of above statement is a trivial extension of the proof of Theorem 3 in (Xu et al., 2019). Briefly, the multiset of the features of the neighbors of node $i$ can be converted from a multiset of tuples of the form $\{\!\!\{(\mu_{ij},h^{(k-1)}_{j}):j\in N(i)\}\!\!\}$ to a multiset of augmented features of the neighbors $\{\!\!\{\tilde{h}^{(k-1)}_{j}:j\in N(i)\}\!\!\}$ where $\tilde{h}^{(k-1)}_{j}=(\mu_{ij},h^{(k-1)}_{j})$. Because $\mu_{ij}$ form a set of finite cardinality, the augmented feature space $\tilde{h}^{(k-1)}_{j}$ is still a countable set and the same proof as in (Xu et al., 2019) can be applied.
@@
Note that $F(i)=\sum_{j\in N(i)}\mu_{ij}f(h^{(k-1)}_{j})$ and $G(i)=\sum_{j\in N(i)}\mu_{ji}f(h^{(k-1)}_{j})$ are only injective up to the sum of the incoming and outgoing weights for a given type of neighbor (see Figure E.1 for an example). This is because the weak form of the WL test only accounts for the sum of the weights of outgoing and incoming edges to neighboring nodes of a given color. This is different from previous work that dealt with undirected graphs (Xu et al., 2019) or graphs with directed but unweighted edges (Rossi et al., 2024).
@@
Because our graphs have a finite number of nodes, $X$ is bounded. For graphs without directed edges or weighted edges, it can be shown (see Lemma 5 of (Xu et al., 2019)) that there always exists an injective function $f$ such that $F(x)=\sum_{x\in X}f(x)$ is injective for all multisets $X$, i.e. if $F(X)=F(Y)$ for two multisets $X$ and $Y$ then $X=Y$. This result can be easily extended to regular directed graphs without weighted edges by considering the neighbors with incoming edges separately from the neighbors with outgoing edges (Rossi et al., 2024). For such graphs, a function $f$ exists such that the tuple $(\sum_{x\in X_{\rightarrow}}f(x),\sum_{x\in X_{\leftarrow}}f(x))$ is only the same for two nodes if they have identical neighborhoods. $X_{\rightarrow}$ and $X_{\leftarrow}$ denotes the multiset of features of the neighboring nodes that are connected to $i$ using an outgoing edge and incoming edge respectively.
@@ F.1 Magnetic Laplacian
$\mathbf{H}^{(q)}=S\odot\exp(\mathrm{i}\Theta^{(q)}).$
$\mathbf{H}^{(q)}$ has some useful properties in capturing the directionality of the edges of the graph. For example, for $q=1/4$, if there is an edge connecting $j$ to $k$ but no edge connecting $k$ to $j$ then $\mathbf{H}^{(q)}_{jk}=\mathrm{i}/2$ and $\mathbf{H}^{(q)}_{kj}=-\mathrm{i}/2$. Although this encoding of edge direction is useful, the fact that both incoming and outgoing edges are purely imaginary (only different up to a sign) means that in general it is impossible to distinguish features aggregated from neighbors that are connected to a node through outgoing edges from those connected through incoming edges. The fuzzy Laplacian $\textbf{L}_{F}$, however, trivially distinguishes features from outgoing neighbors from those of incoming neighbors by keeping one set real and the other imaginary. We will expand on the implication of this for the expressivity of graph neural networks constructed using these two approaches below.
@@ F.2 Limitations of the magnetic Laplacian
Lemma. To prove this statement, we will show that the MLP can learn a linear combination of its input that would it make it equivalent to the graph neural network defined in Eq. 11. This is because when $q=1/8$, the real and imaginary parts of the aggregated features of the neighboring nodes are linear combinations of the features of the outgoing and incoming neighbors.
@@
To extend the magnetic Laplacian to directed graphs with fuzzy edges, we can assign a different value for $q$ (Eq. 15) to each edge. To keep the notation comparable to that of the fuzzy Laplacian, we instead assign a separate angle $\theta_{ij}$ to each edge and define the magnetic Laplacian as, $(\mathbf{L}_{M})_{ij}=e^{\mathrm{i}\theta_{ij}}$ for $j>i$ if there is an edge between the nodes $i$ and $j$, $S_{ij}\neq 0$. To ensure that magnetic Laplacian remains Hermitian, $\mathbf{L}_{M}^{*}=\mathbf{L}_{M}$, we require that $(\mathbf{L}_{M})_{ji}=e^{-\mathrm{i}\theta_{ij}}$, which defines the $\theta_{ij}$ values for $j<i$. It follows that $\theta_{ji}=-\theta_{ij}$.
$\mathbf{L}_{M}=\begin{pmatrix}0&\cdots&\cdots\\ \vdots&\ddots&e^{\mathrm{i}\theta_{ij}}\\ \vdots&e^{-\mathrm{i}\theta_{ij}}&\ddots\\ \end{pmatrix}$
@@
Theorem. For directed graphs with fuzzy edges, the graph neural network constructed using the magnetic Laplacian $\mathbf{L}_{M}$, Eq. 16, is not as expressive as the weak form of the WL graph isomorphism test, and therefore not as expressive as the graph neural network defined using the fuzzy Laplacian $\mathbf{L}_{F}$, Eq. 11.
Proof. The Fuzzy Laplacian matrix conveniently captures the outgoing and incoming weights of the edges from one node to another, $(\mathbf{L}_{F})_{ij}=\cos(\theta_{ij})+\mathrm{i}\sin(\theta_{ij})$. Namely, $\Re(\mathbf{L}_{F})_{ij}=\cos(\theta_{ij})$ is the outgoing weight $\mu_{ij}$ of node $i$ to node $j$ and $\Im(\mathbf{L}_{F})_{ij}=\sin(\theta_{ij})$ is in the incoming weight from node $j$ to node $i$, $\mu_{ji}$. Importantly, the outgoing weight from node $i$ to node $j$, $\mu_{ij}$, is the same as the incoming weight from node $i$ to node $j$. Similarly, the incoming weight from node $j$ to node $i$, $\mu_{ji}$ is the same as the outgoing weight from node $j$ to node $i$. These relationships are directly captured in the fuzzy Laplacian because $(\mathbf{L}_{F})_{ji}=\sin(\theta_{ij})+\mathrm{i}\cos(\theta_{ij})$.
@@
From our definition above, $(\mathbf{L}_{M})_{ij}=\cos(\theta_{ij})+\mathrm{i}\sin(\theta_{ij})$ and $(\mathbf{L}_{M})_{ji}=\cos(\theta_{ij})-\mathrm{i}\sin(\theta_{ij})$. To relate $\theta_{ij}$ to $\mu_{ij}$ we note that the ratio of outgoing and incoming weights for node $i$ is $\mu_{ij}/\mu_{ji}$ and for node $j$ is $\mu_{ji}/\mu_{ij}$. Because these two quantities are reciprocals of each other, we can define